# A 106K Multi-Topic Multilingual Conversational User Dataset with Emoticons

### v1

@@ 3.5.3. Other Sensitive Information
To prevent the leakage of sensitive user information such as names, ages, addresses, organizations, or phone numbers, we use predefined context dictionaries to search the conversation logs. These logs are then further tagged by Llama-3.1-8B-Instruct (Touvron et al., 2023). Any sensitive information found is replaced with the label #SENSITIVE-INFORMATION within the conversation, but are not replaced entirely.
@@ 3.6. Manual Verification
After implementing the automatic process unsafe text, image and privatization we conduct a manual review as an additional safeguard to ensure the dataset’s safety and quality. We perform this in three phases (1) browsing through the conversation for private information leakage and, (2) relabeling them with appropriate tags.
3.7. Conversation Topic Labeling
@@
  • Media Sharing: Conversations in this category often focus on sharing and discussing media content such as emoticons, memes, gifs, and videos for entertainment and social interaction.
@@ 4. Dataset Analysis
We analyze our dataset and share our findings in the following section. Firstly, we present a statistical overview of the dataset. Secondly, we dive into the multi-domain characteristics of our dataset. Finally, we analyze potential user behavior found in our dataset.
@@ 4.1.1. Overall Information
Table 2. Statistics of the raw data and final emoticon dataset. Ours is the largest in size and contains user IDs across conversations.
Field Raw Data Ours
@@
# Avg. Emoticons per User 21.01 16.90
As shown in Table 2, our dataset is large in size and contains user information for cross conversation and cross domain analysis and modeling. Notably, we filter the raw data from the public platform for safety and privacy concerns to form the final dataset.
Table 3. Domain Statistics.
Category Text Messages Emoticons Users Sparsity Dom. Language
Global Unique Avg. Per User Global Use Emoticons
Anime 364,543 87,833 22,572 44.90 10,448 1,956 0.001989 French
Arts 19,515 219 176 3.37 328 65 0.019143 English
Fan Club 551,409 2,287 1,234 13.61 6,341 168 0.011032 Turkish
Finance 2,013,689 45,655 10,083 5.78 245,020 7,894 0.000574 Chinese
Game 2,289,100 79,367 16,382 32.16 48,297 2,468 0.001963 English
Language 2,388,938 51,556 15,801 14.92 50,734 3,456 0.000944 German
Outdoor 353,789 2,088 953 7.88 6,230 265 0.008268 English
Social 416,469 13,077 3,894 11.12 42,180 1,176 0.002856 English
Media Sharing 176,581 87,827 45,451 11.71 14,309 7,498 0.000258 English
Technology 20,305 313 210 4.17 4,425 75 0.019873 English
4.1.2. Emoticon Usage Distribution
Figure 6. Our Dataset Emoticon Usage Distribution. Number of users versus the number of emoticons. Our dataset emoticon usage follows a tail-shaped distribution.
Next, we analyze the distribution of emoticon usages. To better visualize this distribution and prevent distortion, we apply a logarithmic transformation to the number of users. As demonstrated in Figure  6, Our dataset exhibits a tail-shaped distribution, highlighting the real-world emoticon usage patterns.
@@ 4.1.3. Language Distribution
Then, we analyze the language distribution of our dataset, shown in Figure  7. Our dataset features Arabic, Bulgarian, Chinese, Dutch, English, French, German, Hindi, Italian, Japanese, Modern Greek, Polish, Portuguese, Russian, Spanish, Swahili, Thai, Turkish, Urdu, and Vietnamese. English is the most dominant language with an domineering 62.6%. Russian (9.6%) is the second most popular language followed by French (8.6%) and Chinese (8.0%).
Figure 7. Dataset Language Distribution. Our dataset features 20 languages.
4.2. Multi-Topic Characteristics
In this section, we analyze the multi-domain characteristics of our dataset. Firstly, we analyze the domain-group distribution. Secondly, we introduce the domain statistics. Lastly, we analyze the user-emoticon usage overlap across domains.
@@ 4.2.1. Group Distribution
As mentioned above, our dataset contains 67 groups and 10 domains, with the median group being five. We illustrate this in Figure 8. The minimum group count is one, being Arts, and the maximum group count is media sharing. This is due to the nature of resource, focusing on emoticons. Games and Social come close as they tend to be relaxed and so more expressed.
Figure 8. Our dataset features 67 groups and 10 domains. The median group count per domain is five.
@@ 4.2.2. Domain Statistics
Table 3 shows the domain overview statistics. Anime has the most emoticons, followed by Media Sharing and Game, reflecting the expressive nature of these groups. Technology is the sparsest domain, while Media Sharing is the densest, with Finance and Language being unexpectedly dense as well. The average emoticons per user are highest in Anime (44.90) and Game (32.16), while Technology (4.17) and Finance (5.78) have the lowest, suggesting that the nature of the domain influences emoticon usage.
4.2.3. User Emoticon Usage Overlap Across Domains
@@
Figure 9. Distribution of users’ emoticon usage.
Firstly, we establish that users do join other groups and participate; these could be within the same domain or an entirely different domain. The number beside the label on the horizontal axis reveals the total number of emoticon users within the domain, the blue bar represents the percentage of users that appear in other groups within the same domain and the red represents the percentage of users that appear in domains entirely different. We omit domains with fewer than 50 user emoticons to avoid generalization.
Secondly, two of the largest groups with more than 7000 emoticon users (Finance and Media Sharing domain) with roughly 15% and 3% to 85% and 97% different to same domain ratios. In fact, other than the Social domain, domains all have more users participating in the same domain than somewhere else; which could possibly reveal user preference and taste. Naturally, Social domain is clearly an explorative domain and therefore behaves as such.
@@ 5. Application Case 1: User Behavior Analysis
Next, we move on to the user behavior analysis on the dataset. More specifically, we demonstrate our dataset’s potential on showcasing (1) user distinct style (2) multi-domain behavior (3) temporal behavior change (4) user behavior with different respondents.
@@ 5.1. Distinct User Styles
Figure 10. User A emoticons in the Language group.
Figure 11. User B emoticons in the Game group.
As shown in Figure 10, User A’s emoticon styles are predominantly cat-themed and cute, while User B’s emoticons are more idol-based, with an emphasis on adorable and charming designs, as shown in Figure 11. This demonstrates that users generally have a distinct emoticon preference and this potentially be used for user modeling.
@@ 5.2. Cross Domain Behavior Change
More interestingly, we analyze if user, due to different context, could change their style. Herein, the same User A, Figure 12, displays a significantly different style in the Media Sharing group, which is more structured and cartoonish in appearance. This is contrastive to the originally cartoonish style.
Figure 12. User A emoticons in the Media Sharing group.
Figure 13. User B emoticons in the Outdoor group.
On the other hand, User B’s emoticons, Figure 13, still retain elements of their signature adorable style but are subtly tuned to match the outdoor-themed context of the new domain, which is about rats. This behavior reveals that different user have different consistency level in emoticon-usage. There are possibly many factors for this behavior such as the group context.
@@ 5.3. Temporal Behavior Change
Figure 14. User C emoticon usage over time where count is more than three.
We collate the top emoticon by usage count from User C and present in Figure  14. As can be seen, the user’s emoticon preferences evolve over time. In September and October, the user gravitates toward dog-themed, cartoonish emoticons. However, by January, their preference shifts to smaller, cuter dog emoticons. Interestingly, in June, the user shows a renewed interest in a different set of realistic dog emoticons, while still maintaining a fondness for the smaller, cuter dog designs.
Overall, despite fluctuations in style, the user consistently displays a preference for dog emoticons, with a notable emphasis on the cute, cartoonish variations. This suggests that while the user’s style may vary, their core preference remains for dog-themed emoticons, particularly those with a cute and cartoonish aesthetic.
Figure 15. User D emoticon usage over time. The user uses the green frog emoticon frequently overtime.
On the other hand, User D demonstrates a consistent use of a certain green laughing frog, except the months where no emoticons are used, as seen in Figure  15. Herein, the user demonstrates a consistent usage of a certain emoticon, unlike User C.
5.4. User Behavior with Different Respondents
Figure 16. User E shows a consistent style preference when replying to Users F and G.
We present two distinct cases of user behavior in response to different users. In the first case, User E replies to both Users F and G using a similar set of emoticons. As shown in Figure  16, User E consistently selects a boy emoticon with a nonchalant expression, regardless of the recipient. Although the third and fourth emoticons in the sequence differ, the first two emoticons remain frequently used, with counts of 12 and 8 respectively, indicating a strong preference for these emoticons.
@@ 5.4. User Behavior with Different Respondents
Figure 17. User H shows contrasting style preferences when replying to Users I and J.
In contrast, User H exhibits a clear difference in emoticon usage when replying to Users I and J. As illustrated in Figure  17, User H uses bubbly, cheerful, and generally cute emoticons when replying to User I. However, when replying to User J, User H switches to a more melancholic and indifferent style, favoring stickman emoticons that are notably less cute. This shift in emoticon preference highlights the variation in user behavior based on the recipient of the message.
In conclusion, these examples demonstrate that the dataset captures meaningful differences in emoticon usage across various contexts. This variation, influenced by the person being replied to, suggests the potential for deeper personalization and more insightful user behavior analysis.
6. Application Case 2: Emoticon Recommendation
@@
In this section, we evaluate the feasibility of our dataset through quantitative experiments. Specifically, we evaluate our dataset on current emoticon recommendation.
@@ 6.1. Experiment Settings
We evaluate two baseline methods for emoticon recommendation, described as follows:
@@
We conduct experiments on three distinct perspectives of the dataset: (1) the English and Chinese subsets, (2) the domain-specific subset, and (3) the complete dataset.
6.1.1. English and Chinese Subset
Since MOD can only handle English and Chinese, we select groups with more than 80% of text being either English or Chinese and train the model on it. We do the same for SRS.
6.1.2. Domain
To interpretate the performance between different domains, we separate the dataset by domains as shown in Table 3. Since SRS is not limited by multilingualism, we evaluate using SRS.
6.1.3. Complete
Finally, the entire dataset is utilized for evaluation on SRS.
@@ 6.2.1. English and Chinese Subset
Table 4. emoticon Retrieval Results for MOD and SRS (English and Chinese). We shortform Negative Samples for NS.
@@ 6.2.2. Domain
Table 5 reveal that the emoticons domain performs the best overall, with an MRR of 0.335, R@1 of 0.143, R@3 of 0.365, and R@5 of 0.552. On the other hand Arts domain has a relatively lower performance, possibly due to the higher sparsity (0.019143).
Table 5. Results of SRS on individual domains.
@@ 6.2.3. Complete Dataset
Table 6. Performance metrics for varying negative samples across multiple evaluation criteria on the entire dataset using SRS method.
@@ 7. Discussions
7.1. Other Potential Applications
Beyond user behavior analysis and emoticon recommendation, the dataset presents several opportunities for further research. One potential avenue is a more extensive User Behavior Modeling, which could involve analyzing temporal changes in user behavior and investigating whether emoticon usage varies based on the responder. Additional potential applications include, but are not limited to, Personalized Emoticon Retrieval and Generative AI, both of which have previously faced dataset limitations (Shen et al., 2024; Xu et al., 2024).
@@ 8. Conclusion
In this paper, we introduce our emoticon dataset, the largest emoticon dataset to date, comprising data from 22K users, 370.2K emoticons, and 8.3M conversation messages. This multi-domain dataset encompasses a wealth of diverse information, capturing temporal, multilingual, and cross-domain behaviors not found in previous datasets. Extensive quantitative and qualitative experiments demonstrate the practical applications of our dataset in user behavior modeling and personalized emoticon recommendation. It also holds potential for further research in areas such as personalized retrieval and conversational studies.

### v2

@@ 3.5.3. Other Sensitive Information
To prevent the leakage of sensitive user information such as names, ages, addresses, organizations, or phone numbers, we use predefined context dictionaries to search the conversation logs. These logs are then further tagged by Llama-1.5b (Liu et al., 2023). Any sensitive information found is replaced with the label #SENSITIVE-INFORMATION within the conversation, but are not replaced entirely.
@@ 3.6. Manual Verification
After implementing the automatic process unsafe text, image and privatization we conduct a manual review as an additional safeguard to ensure the dataset’s safety and quality. In this phase, a subset of the dataset is randomly sampled and carefully examined by human reviewers. The goal is to verify that the automated procedures have successfully removed or replaced all sensitive and harmful content without affecting the overall structure or meaning of the messages. The manual review process is essential for catching any potential issues that automated processes may have missed, such as false negatives or inaccuracies in labeling.
Finally, 370,222 (72.28%) of 512,192 stickers do not contain offensive taggings, 8,286,422 (99.45%) of 8,332,351 utterances are safe and private. Then, there are a remaining 67 conversation groups.
3.7. Conversation Domain Labeling
@@
  • Media Sharing: Conversations in this category often focus on sharing and discussing media content such as stickers, memes, gifs, and videos for entertainment and social interaction.
@@ 4. Dataset Analysis
We analyze the U-Sticker dataset and share our findings in the following section. Firstly, we present a statistical overview of the dataset. Secondly, we dive into the multi-domain characteristics of U-Sticker. Finally, we analyze potential user behavior found in U-Sticker.
@@ 4.1.1. Overall Information
Table 2. Statistics and Comparison of Raw Data and U-Sticker (A) and Top-8 Languages Post-Processing (B).
A. Statistics Comparison.
Field Raw Data U-Sticker
@@
Avg. Stickers/User 20.54 16.90
B. Languages.
Top-8 Languages
1. English (62.6%)
2. Russian (9.6%)
3. French (8.6%)
4. Chinese (8.0%)
5. Spanish (2.5%)
6. Turkish (2.0%)
7. Vietnamese (1.9%)
8. Polish (1.5%)
As shown in Table 2, our U-Sticker is large in size and contains user information for cross conversation and cross domain analysis and modeling. Notably, we filter the raw data from the public platform for safety and privacy concerns to form the final U-Sticker dataset.
Table 3. Domain Statistics. Notably, Anime, Media Sharing, and Game are the most active domains in terms of sticker usage, while Technology is the least active. The sparsity values highlight the varying levels of engagement across domain. Sparsity values are scaled by a factor of 10,000 for easier readability.
Category Sticker Users Sparsity
Global Unique Global Sticker
Anime (4) 87,833 22,572 10,448 1,956 19.89
Arts (1) 219 176 3.37 65 191.43
Fan Club (4) 2,287 1,234 6,341 168 110.32
Finance (8) 45,655 10,083 245,020 7,894 5.74
Game (1) 79,367 16,382 48,297 2,468 19.63
Language (5) 51,556 15,801 50,734 3,456 9.44
Outdoor (5) 2,088 953 6,230 265 82.68
Social (9) 13,077 3,894 42,180 1,176 28.56
Media (18) 87,827 45,451 14,309 7,498 2.58
Tech (2) 313 210 4,425 75 198.73
4.1.2. Sticker Usage Distribution
Figure 3. Sticker Usage Distribution. Number of users against the number of stickers. U-Sticker usage trend follows a long-tail distribution, highlighting the real-world sticker usage patterns.
Next, we analyze the distribution of sticker usages. To better visualize this distribution and prevent distortion, we apply a logarithmic transformation to the number of users. As demonstrated in Figure  3, the U-Sticker usage trend follows a long-tail distribution, where most users employ between 0 and 100 stickers, while only a small fraction exhibit extremely high usage. This behavior is consistent with Zipfian or Pareto-like distributions (Zipf, 1949) (new, [n. d.]), and is typical of organic user activity observed in large-scale, real-world datasets.s.
@@ 4.1.3. Language Distribution
Then, we analyze the language distribution of U-Sticker, shown in Figure  2. U-Sticker features Arabic, Bulgarian, Chinese, Dutch, English, French, German, Hindi, Italian, Japanese, Modern Greek, Polish, Portuguese, Russian, Spanish, Swahili, Thai, Turkish, Urdu, and Vietnamese. English is the most dominant language with an domineering 62.6%. Russian (9.6%) is the second most popular language followed by French (8.6%) and Chinese (8.0%).
4.2. Multi-Domain Characteristics
In this section, we analyze the multi-domain characteristics of U-Sticker. Firstly, we analyze the domain-group distribution. Secondly, we introduce the domain statistics. Lastly, we analyze the user-sticker usage overlap across domains.
@@ 4.2.1. Group Distribution
As mentioned above, U-Sticker contains 67 groups and 10 domains, with the median group being five. The minimum group count is one, being Arts, and the maximum group count is media sharing. This is due to the nature of resource, focusing on stickers. Games and Social come close as they tend to be relaxed and so more expressed.
@@ 4.2.2. Domain Statistics
We present the overview statistics of domain in Table 3. Looking at sticker global counts, the domain with the most sticker is Anime, followed by Media Sharing and Game. This is reasonable as these are highly relaxed groups and tend to be more expressive, thereby having a higher sticker usage count.
The most sparse domain is Technology and the most dense domain is Media Sharing. This is natural as technology is a rather technical topic and audience could be less inclined to using sticker expressions. On the other hand, media sharing sometimes encourage chats entirely comprising of stickers, hence increasing the density. Surprisingly, Finance and Language are also denser domains. This could be because Finance and Language are semi-professional scenarios which could call for an increased sticker usage.
Lastly, the average stickers used per user varies largely with Anime and then game domain having the highest 44.90 and 32.16 respectively. On the other hand, Technology and Finance domains have 4.17 and 5.78 stickers respectively. We hypothesize that the major difference in the nature of the scenario affects sticker usage in users.
4.2.3. User Sticker Usage Overlap Across Domains
@@
Figure 4. Distribution of users’ sticker usage. Social domain users features the largest distribution of users in contrastive (yellow), reflecting the inherently explorative nature of their domain.
Firstly, we establish that users do join other groups and participate; these could be within the same domain or an entirely different domain. The number beside the label on the horizontal axis reveals the total number of sticker users within the domain, the blue bar represents the percentage of users that appear in other groups within the same domain. We omit domains with fewer than 50 user stickers to avoid generalization.
Secondly, two of the largest groups with more than 7000 sticker users (Finance and Media Sharing domain) with roughly 15% and 3% to 85% and 97% different to same domain ratios. In fact, other than the Social domain, domains all have more users participating in the same domain than somewhere else; which could possibly reveal user preference and taste. Naturally, Social domain is clearly an explorative domain and therefore behaves as such.
@@ 5. Application Case 1: User Behavior Analysis
Next, we move on to the user behavior analysis on the U-Sticker dataset. More specifically, we demonstrate U-Sticker’s potential on showcasing (1) user distinct style (2) multi-domain behavior.
@@ 5.1. Distinct User Styles
As shown in Figure 5, User A’s sticker styles are predominantly cat-themed and cute, while User B’s stickers are more idol-based, with an emphasis on adorable and charming designs. This demonstrates that users generally have a distinct sticker preference and this potentially be used for user modeling.
Figure 5. User A stickers in the Language group and User B stickers in the Game group.
@@ 5.2. Cross Domain Behavior Change
Figure 6. User A stickers in the Media Sharing group and User B stickers in the Outdoor group.
More interestingly, we analyze if user, due to different context, could change their style. Herein, the same User A, Figure 6, displays a significantly different style in the Media Sharing group, which is more structured and cartoonish in appearance, vastly contrastive to the originally cartoonish style. On the other hand, User B’s stickers, still retain elements of their signature adorable style but are subtly tuned to match the outdoor-themed context of the new domain, which is about rats. This behavior reveals that different user have different consistency level in sticker-usage. There are possibly many factors for this behavior such as the group context.
@@ 5.3. Temporal Behavior Change
We present the top-used stickers of User C in Figure  7, showing how their preferences shift over time. In September and October, they favor cartoonish dog stickers; by January, they prefer smaller, cuter dog designs; and in June, they lean toward realistic dog stickers while still using the cute ones. Despite style changes, their consistent preference for dog-themed stickers—especially cute ones—remains.
Figure 7. User C’s evolving sticker preferences over time, shifting from expressive dog stickers in September to smaller, more reserved dog stickers by January, with a renewed interest in realistic dog stickers by June, maintaining a consistent preference for dog-related designs.
In contrast, User D, Figure  8 consistently uses the same green laughing frog sticker (commonly known as ”Pepe”) across many months, with few gaps. Overall, this demonstrates that our dataset contains both types of users: those like User D, who exhibit stable preferences for specific stickers, and those like User C, whose preferences evolve over time. Intuitively, this aligns with common sense, as people can have both consistent and fluctuating preferences depending on various factors. This diversity in our dataset enhances the accuracy and realism of the personalized sticker retrieval task, allowing it to better reflect the range of behaviors found in real-life scenarios.
Figure 8. User D’s consistent usage of the green laughing frog sticker (Pepe frog) from February to January, with occasional months of no sticker usage, highlighting a strong preference for this particular sticker over time.
5.4. User Behavior with Different Responders
We highlight two examples of how sticker usage varies with different recipients. In the first case, Figure 9, User E replies to both Users F and G using a consistent set of stickers—primarily a nonchalant boy sticker—demonstrating stable preferences regardless of recipient.
Figure 9. User E consistently uses the red t-shirt with smirk boy sticker and the grey cartoon straight-face sticker when replying to Users F and G
@@ 5.4. User Behavior with Different Responders
Figure 10. User H displays contrasting style preferences when replying to Users I and J, using cute, cuddly bear stickers for User I and black-and-white, melancholic stickers for User J.
Contrastively, User H adapts their sticker style depending on the recipient: using cheerful, colorful stickers with User I, but switching to melancholic, monochromatic stickers with User J, Figure 10.
In conclusion, these examples demonstrate that our dataset captures meaningful variations in sticker usage based on the context of the interaction. This variability, driven by the recipient, intuitively aligns with our commonsense understanding of how individuals communicate differently with different people. U-Sticker effectively captures this behavior, which in turn empowers related tasks, such as personalized sticker retrieval, by enabling a deeper and more context-aware analysis of user preferences.
6. Application Case 2: Sticker Recommendation
@@
In this section, we evaluate the feasibility of our dataset through quantitative experiments. Specifically, we evaluate U-Sticker dataset on current sticker recommendation.
@@ 6.1. Experiment Settings
We evaluate two baseline methods for sticker recommendation, described as follows:
@@
We conduct experiments on three distinct perspectives of the U-Sticker dataset: (1) the English and Chinese subsets, (2) the domain-specific subset, and (3) the complete dataset.
Since MOD can only handle English and Chinese, we select groups with more than 80% of text being either English or Chinese and train the model on it. We do the same for SRS. This forms the English and Chinese Subset setting. To interpretate the performance between different domains, we separate the dataset by domains as shown in Table 3. Since SRS is not limited by multilingualism, we evaluate using SRS. This forms the Domain setting. Finally, the complete U-Sticker dataset is evaluated on SRS.
@@ 6.2.1. English and Chinese Subset
Table 4. Sticker Retrieval Results for MOD and SRS on the English and Chinese. Negative Samples is shortformed NS.
@@ 6.2.2. Domain
Table 5 reveal that the Media Sharing domain performs the best overall, with an MRR of 0.335, R@1 of 0.143, R@3 of 0.365, and R@5 of 0.552. On the other hand Arts domain has a relatively lower performance, possibly due to the higher sparsity (0.019143).
Table 5. Results of SRS on individual domains. Media Sharing domain performs the best overall, whereas Arts domain has a relatively lower performance possibly due to higher sparsity.
@@ 6.2.3. Complete Dataset
Table 6. Performance metrics for varying NS across multiple evaluation criteria on the entire dataset using SRS method.
@@ 7. Discussions
7.1. Other Potential Applications of U-Sticker
Beyond user behavior analysis and sticker recommendation, the U-Sticker dataset presents several opportunities for further research. One potential avenue is a more extensive User Behavior Modeling, which could involve more quantitative user analysis methods, or involve analyzing temporal changes in user behavior and investigating whether sticker usage varies based on the responder. Additional potential applications include, but are not limited to, Personalized Sticker Retrieval and Generative AI, both of which have previously faced dataset limitations (Shen et al., 2024; Xu et al., 2024).
@@ 8. Conclusion
In this paper, we introduce U-Sticker dataset, the largest sticker dataset to date, comprising data from 22.6k users, 370.2k stickers, and 8.8M conversation messages. This multi-domain dataset encompasses a wealth of diverse information, capturing temporal, multilingual, and cross-domain behaviors not found in previous datasets. Extensive quantitative and qualitative experiments demonstrate the practical applications of U-Sticker in user behavior modeling and personalized sticker recommendation. It also holds potential for further research in areas such as personalized retrieval and conversational studies.